Kusho Industries produces and sells computer chips. Its (hourly) production function is 𝒒 = πŸ’π‘²πŸŽ.πŸ’π‘³πŸŽ.πŸ”, while its (hourly) cost function is 𝒄 = πŸπŸŽπ‘³ + πŸ–πŸŽπ‘². Furthermore, Kusho must produce π’’πŸŽ = πŸ’πŸŽπŸŽ computer chips per hour.

a. Which levels of 𝑳 and 𝑲 satisfy the first-order conditions for the constrained minimisation of Kusho’s

cost? Use the Lagrange Multiplier (LM) method. Also, find and interpret the value of the Lagrange

multiplier (𝝀). [8]

b. Show that 𝑴𝑹𝑻𝑺 = π’˜ at the constrained cost minimising levels of 𝑳 and 𝑲 obtained above. [2]


1
Expert's answer
2021-06-22T17:47:42-0400

As per Lagrange method, the cost is minimized subject to production function.

Β Minimize the"C = 20L + 80K" subject to "Q = 4 K^{0.4}L^{0.6}"

Lagrange Function:

"R= 20L + 80K - \u03bb(4 K^{0.4}L^{0.6} \u2013 Q)"

Maximization of R with respect to L and K,

"R_L = 20 – λ (4 \times 0.6 L{-0.4}K^{0.4})"

"R_L = 0"

"λ = (\frac{25}{3}) (\frac{L}{K})^{0.4} ............(1)"

"R_K = 80 – λ (4 \times 0.4 L^{0.6} K{-0.6})"

"R_K = 0\\

λ = 50 (\frac{K}{L})^{0.6}..................(2)\\

R λ = - λ (4 K^{0.4}L^{0.6} – Q)\\

R λ = 0\\

Q = 4 K^{0.4}L^{0.6} ....................(3)\\"

Equating both value of Ξ» as follows:

"L = 6K ..................(4)"

Putting equation (4) in equation (3) as follows:

"Q = 4 K^{0.4}L^{0.6}\\

400 = 4 K^{0.4}L^{0.6}\\

400 = 4 K^{0.4}(6K)^{0.6}\\

K = 34.12\\

L = 6 \times 34.12\\

= 204.76"

The value of Lagrange multiplier is calculated as follows:

"λ = 50 (\frac{K}{L})^{0.6}\\ 

= 50 \times 0.3412\\

= 17.06"

It shows the change in output due to one unit change in labor or capital.


b.

The MRTS is calculated as follows:

"MRTS = \frac{MP_L}{MP_K}\\

= \frac{3K}{2L}\\

= 3 \times \frac{34.12}{2}\times204.76\\

= 0.25"

The value of MRTS will be equal to ratio of wage rate and rental rate.

"\frac{w}{r} = \frac{20}{80}\\

= 0.25"

At equilibrium, the MRTS is equal to ratio of wage rate and rental rate.